# Moderate Nesting and Cross-Equatorial Asymmetry of Active Regions in Solar Cycle 24

### v1

@@ Abstract
Solar Cycle 24 data are used to determine how often the Sun emerges sunspots in ‘activity nests’, i.e., regions where sunspots and active regions (ARs) repeatedly emerge. We use the Solar Photospheric Ephemeral Active Region (SPEAR) catalog created from Helioseismic and Magnetic Imager (HMI) data as well as the HMI Carrington Rotation maps of radial magnetic field, . The Sun shows moderate nesting behavior with 41% (48%) of AR magnetic flux found in Northern (Southern) hemispheric nests that are short-lived (4 months). Different rotation rates are used to search for nests that may not be evident ‘by eye’. The maximum number of nests are found with slightly prograde rotational velocities, with significant nest flux also found at synodic 451–452 nHz prograde and 409–411 nHz retrograde frequencies. Nest patterns show strong hemispheric asymmetry, indicating that the physical origin of nests identified herein must also be asymmetric or antisymmetric across the equator.
@@ 1 Introduction
Magnetism in Sun-like and lower-mass stars is ubiquitous and dynamic, shaping their immediate space environments. These magnetic fields arise from a dynamo process operating in the stellar interior that amplifies and organizes magnetic energy. Magnetic flux emerges into the solar atmosphere in the form of dark sunspots, or active regions (ARs), which suppress convection and locally alter the transport of energy.
While ARs have been observed for centuries, the physical processes that give rise to them remain an open question. In particular, it is not fully understood whether the flux that produces an AR is generated at the base of the convection zone near the tachocline or within the bulk of the convection zone itself. Although it is widely accepted that magnetic flux tubes rise through the convection zone via buoyancy, the manner in which flux organizes and emerges into the atmosphere remains poorly constrained (Fan2009; Cheung2010).
A striking observational feature of ARs is their tendency to cluster in longitude and persist in certain regions of the Sun for extended periods. These so-called activity nests are regions where multiple ARs emerge within a limited longitude–latitude range over a span of several Carrington rotations. However, the solar community lacks an agreed-upon quantitative definition of what constitutes an activity nest. Nests have been identified using photospheric sunspot catalogs, magnetograms, and coronal emission maps, with each data source providing slightly different perspectives on clustering.
The term activity nest is often used synonymously with active longitude. Here, however, we propose that they describe different behaviors. We use activity nests to describe the clustering of ARs in one hemisphere, while we use active longitudes to describe the tendency of ARs to emerge at similar longitudes in opposite hemispheres, i.e., with symmetry across the equator, see Table 1. The term hot spot is used to describe regions of solar activity with a high concentration of flares and sunspot groups. The longitudinal distribution of solar magnetic activity is therefore a probe of the non-axisymmetric components of the solar dynamo, which remain relatively unexplored.
Castenmilleretal1986 proposed a quantitative definition: at least three ARs emerging within four Carrington rotations, within 7.5∘ longitude and 5∘ latitude of one another. They find that 30-40% of ARs in the years 1959-1964 participated in nests. Earlier studies also highlighted persistent longitudes that hosted activity, sometimes termed “active longitudes” (Bogart1982; Bai1987; Bai2003; Henney2002). berdyugina:2003 analyzed long-term sunspot records and suggested the existence of two nests in a given hemisphere, roughly 180∘ apart, persisting for centuries with the dominance alternating between the nests on the order of 3 years — a phenomenon described as flip-flop. Gyenge2014; Gyenge2016; Gyenge2017 investigated the temporal evolution of active longitudes across solar cycles, while Usoskin2005; Usoskin2007 analyzed flip-flop patterns in long-term sunspot series. The identification of long-lived active longitudes relies on analysis techniques that often select only the two most prominent active regions in any given Carrington rotation, and fit this data with bolometric curves. The analysis techniques were criticized by pelt:2005 who showed that random input data can result in similar longitudinal distributions when subjected to similar filtering and analysis. mandal:2017 revisited this type of analysis using digitized sunspot data from the Kokaikanal Observatory to confirm long-lived active longitudes with a 180∘ separation. More recently, Finley2024 identified a persistent nest in Cycle 25 observed in SDO/AIA, noting its role in anchoring the heliospheric current sheet.
| Term | Usage and References |
@@
| enhanced flaring; Bai (2003) |
Figure 1 depicts cartoon distributions of sunspots: (left) moderate nesting similar to the Sun, (middle) moderate nesting with sunspots aligned along active longitudes, and (right) strong nesting as might be observed on a sun-like star, combined with well-defined active longitudes.
The physical origins of activity nests remain uncertain. Possible mechanisms include the influence of giant convective cells, magneto-Rossby waves, or MHD instabilities acting on toroidal flux bands near the tachocline (Zaqarashvili2010; Dikpati2018). Flux emergence simulations also suggest that clustering may reflect subsurface connectivity of buoyant magnetic structures (Fan2009; Cheung2010). Persistent longitudinal preferences may indicate non-axisymmetric dynamo modes originating primarily from global MHD instabilities (chan2004; bigazzi2004) or interactions between global rotation and convection.
The identification of activity nests, hot spots, or active longitudes is sensitive to the rotation rate assumed in the analysis. Most studies have searched for preferred longitudes in a reference frame rotating close to the Carrington rate, implicitly assuming that longitudinal structures remain anchored in that frame (e.g., Bai1987; Bai2003). However, different choices of rotation rate can highlight or obscure persistent features. For example, Bogart1982 applied Fourier analysis to sunspot numbers and found prominent rotation periods that varied by as much as two days between solar cycles, suggesting that active longitudes may not be rigidly tied to a single rate. Likewise, Berdyugina2003; Usoskin2005 reported that active longitudes can migrate with differential rotation, exhibiting variable rotation rates that depend on both latitude and phase of the solar cycle. Analyses of hot spots have often assumed a rigid rotation over decades, which can impose artificial persistence if the underlying pattern is, in fact, drifting. In this work, we expand on these approaches by searching across a broad range of rotation rates, allowing us to identify nests and longitudinal clustering without presupposing their anchoring frame.
The Sun provides a near-by star in which to study these phenomena in the broader context of stellar activity. Long-lived active longitudes have been reported on other stars using data from Kepler and TESS (e.g. Lehtinen2016; Lehtinen2020; Reinhold2020; Davenport2015; Roettenbacher2016). Dynamo models also suggest that non-axisymmetric modes and flux-emergence asymmetries may be generic features of rapidly rotating stars (Isik2011; Isik2015; Brun2017). Understanding solar activity nests and active longitudes therefore informs not only solar-cycle physics but also stellar magnetism and the resulting brightness variations that impact exoplanet detection and habitability studies (Baliunas1995).
In summary, activity nests and active longitudes represent key diagnostics of longitudinal organization in solar and stellar magnetism. Persistent clustering implies memory in the dynamo system, or long-lived dynamic structuring, that is not captured by purely axisymmetric flux-transport models.
@@ 2 Data
For this study, we employ synoptic Carrington rotation maps of the vector magnetic field, , from the Helioseismic and Magnetic Imager (HMI; Scherrer2012; Schou2012) on board the Solar Dynamics Observatory (SDO). The synoptic maps provide full-Sun coverage of the photospheric magnetic field by assembling central meridian observations over a full Carrington rotation of 27.2753 days, enabling consistent analysis of the longitudinal distribution of solar activity. The maps can be accessed through JSOC in the data series hmi.synoptic_mr_polfil_720s. The dimensions are 3600 1440 with equal steps in longitude and sine latitude with values in the units of Mx cm-2. In addition, we make use of the Solar Photospheric Ephemeral Active Region (SPEAR) catalog (http://hmi.stanford.edu/hminuggets/?p=3730), which is based on the Space-weather HMI Active Region Patch (SHARP) data product. SHARPs identify and track ARs in near-real time using HMI vector magnetograms (Bobra2014), and the SPEAR catalog provides a tabulated set of data by compiling AR properties such as total flux, latitude and longitude location, tilt angle, etc., when the SHARP regions are nearest to the central meridian. Each SHARP region is only found once in the catalog.
@@ 3 Methods
We use HMI Carrington Rotation maps of the radial vector magnetic field, , to visually identify the times and locations of strong active region magnetic flux. We plot the absolute value of between the latitudes of , keeping the hemispheres separate in order to note the northern and southern hemisphere patterns. We downsample the Carrington longitude from 3600 to 360 pixels and mildly smooth the signal in latitude and longitude. Since the Sun rotates differentially with higher latitudes rotating slower than those close to the equator, the Carrington rate is only the value used to build the synoptic maps and not indicative of the rate at which magnetic features rotate.
When using the SPEAR catalog to identify nests, we remove long-lived ARs that return to the front side of the Sun for a second or third time. We do this using a method developed by Yeates:2020 using flux transport models to estimate what a given AR rotating off the front side of the disk would look like upon its return. If the forward-modeled AR is a good match for the observed AR, then the observed AR NOAA number and SHARP number is listed as a returning sunspot and removed from being a possible candidate for a member of a nest.
To identify a nest, each AR is used as a starting data point, then nearby latitudes, longitudes and Carrington Rotations are searched to find other ARs that could be part of the nest given a certain rotation rate and nest definition. If a rotation rate is specified that is different from the Carrington rate of 424.3427 nHz, then the search for ARs in time updates the search locations to be longitude corresponding to the change in longitude per rotation that would correspond to that rotation rate. If an AR is identified in two nests, then those nests are merged. All frequencies in this paper are synodic, not sidereal. To convert to sidereal rates, add 31.6875 nHz to the synodic rate.
@@
In order to test the hemispheric asymmetry (i.e., whether ARs or nests tend to occur at the same longitude across the equator or not) and the distribution, both of the active regions in Cycle 24 and the identified nests, we plot locations of ARs and nests as a function of Carrington longitude and Carrington rotation and calculate how often they intersect or occupy the same longitude. We report on how likely it is for an AR or nest to exist in the same longitude and time bin as well as how likely this value is in a randomized distribution. For this test, the longitudes are randomized for each AR but the nest longitudes are randomized for the entire nest (keeping the ARs close in longitude for a given nest). ARs are slightly randomized in time but with a criteria that the total number of ARs in each hemisphere are constant within a given year.
@@ 4 Results
The amount of nesting exhibited by the Sun depends on the definition used to describe what constitutes an activity nest. Early attempts at nest detection were frustrating because the lifetimes and numbers of ARs in each nest were simply reflections of the adopted definition. Therefore, a larger parameter space with varying criteria were used to search for activity nests.
@@ 4.1 Degree of Nesting
Using the criteria of Castenmilleretal1986, in which the nest critera is three ARs found within four CRs with 7.5∘ longitude and 5∘ latitude of each other, the SPEAR catalog is used to search through the AR locations to find clustering. When using the Carrington rate to search the catalog, 40 nests with 25% of ARs and 32% of flux are identified in the North and 30 nests with 23% of ARs and 35% of flux are found in the South, see panel b in Figures 2 and 3 that shows separate nests plotted as different symbols and colors. Note that latitude is not shown in the figures so that even if there is a large amount of flux in a location, a nest might not be found if the ARs are further away from each than 5∘ latitude.
The number of ARs and AR flux in nests increases when the data is searched using slightly prograde rates of rotation. The maximum amount of AR flux participating in nests is 48% (41%) in the Southern (Northern) hemisphere at rotational rates of 428.5 (426.7) nHz, slightly prograde with respect to the Carrington rate, see panel c in Figures 2 and Figure 3. These are considered short-lived nests.
@@ 4.2 Nests at Different Rotation Rates
The percent of flux found in nests as a function of rotation rate is shown in Figure 4 for the Castenmilleretal1986 nest criteria, i.e. 3 ARs in 4 CRs, basically short-lived nests. The North and South hemisphere show significant flux in nests for rates near the Carrington rotation rate, i.e. synodic frequencies of 418-433 nHz synodic, or 27.7 -26.7 days. Other frequencies that contain significant flux in nests are found at 451-452 nHz and 409-413 nHz. These may be related to other processes such as inertial modes.
@@ 4.3 Hemispheric Asymmetry
We examine the locations of all ARs that are assigned a NOAA number and also have an area greater than 50 Hemispheres. Separating the ARs into North and South hemispheres, we plot them as a function of Carrington Rotation and Carrington Longitude, see Figure 5, top panel, with North and South being red and blue symbols. After dilating their locations by longitude and 3 Carrington rotations to provide some width, we count the percentage of ARs that are coincident in time (3 CRs) and longitude () across the Northern and Southern hemisphere. The fraction of ARs that overlap in North and South bins with this dilation is 58%. This can be interpreted as follows: given an AR in the northern hemisphere, there is a 58% chance that an AR will be located within 9∘ longitude and 3 Carrington rotations in the southern hemisphere.
The same is done for the nest population of ARs, see Figure 5, middle panel, using the Castenmiller criteria with a Carrington rotation rate. Out of all of the longitude–time bins occupied by an AR in a nest in either hemisphere, only 8% are found to have an AR in the opposite hemisphere within 3 CRs in time and in longitude, while more than 92% are exclusive to the North or South alone without an AR in the opposite hemisphere. This means that when nests form in one hemisphere, they are overwhelmingly absent at the corresponding longitude and time in the opposite hemisphere. In contrast to the broader active region population, which shows substantial co-occupancy, the organization of nests is characterized by strong hemispheric antisymmetry. Identifying nests using a slightly prograde rotation rate (similar to the rates shown in Figures 2 and 3 when the maximum amount of flux is found in nests), but with a similar dilation of 3 CRs in time and in longitude, provides a North-South AR intersection of 11.9% which is more antisymmetric than 83.4% of trials.
It is difficult to determine the significance of the nest asymmetry by randomizing the nest locations since the existence of nests identified with this criteria are already 3 above the noise. We try to characterize the asymmetry by randomly shifting the average nest locations in longitude (but keeping the number of ARs in the nest constant) and shifting the average time of the nest forward or backward in time with the limitation that the shift is less than a year. The 10,000 trials show that the observed asymmetry is more than 92% of the randomized trials, see Figure 5 lower panel. While this is not above a 2 signal, we believe it is significant due to the fact that we cannot truly randomize the nest locations since searching for nests in AR locations that are randomized in longitude leads to much fewer and often zero nests.

### v2

@@ Abstract
Solar Cycle 24 data are used to determine how often the Sun emerges sunspots in ‘activity nests’, i.e., regions where sunspots and active regions (ARs) repeatedly emerge. We use the Solar Photospheric Ephemeral Active Region (SPEAR) catalog created from Helioseismic and Magnetic Imager (HMI) data as well as the HMI Carrington Rotation maps of radial magnetic field, . The Sun shows moderate nesting behavior with 41% (48%) of AR magnetic flux found in Northern (Southern) hemispheric nests that are short-lived (average lifetimes 3.3 - 4.0 months). Different rotation rates are used to search for nests that may not be evident ‘by eye’. The maximum number of nests are found with slightly prograde rotational velocities, with significant nest flux also found at synodic 451–452 nHz prograde and 409–411 nHz retrograde frequencies. Nest patterns show strong hemispheric asymmetry, indicating that the physical origin of nests identified herein must also be asymmetric or antisymmetric across the equator.
@@ 1 Introduction
Cool stars (0.2–1.3 ) like the Sun generate dynamic magnetic fields that influence planetary environments. These fields arise from a dynamo operating in the stellar interior that amplifies and organizes magnetic fields (Parker, 1975). Magnetic flux emerges into the solar atmosphere in the form of sunspots, or active regions (ARs), that are magnetic in nature (Hale, 1908) and suppress convection and locally alter energy transport.
Although ARs have been observed for centuries, the origin and organization of the magnetic flux that produces them remain open questions. Magnetic flux emergence is widely attributed to the buoyant rise of magnetic structures through the convection zone, yet the depth and organization of the source field remain uncertain (Fan, 2009; Cheung et al., 2010).
A striking observational feature of ARs is their tendency to cluster in longitude and persist in preferred regions of the Sun for extended periods. These regions, often termed activity nests, are characterized by multiple ARs emerging within a limited longitude–latitude range over several Carrington rotations. This was noted as early as 1863 by Carrington who suspected that active regions were not uniformly and randomly distributed in longitude (Carrington, 1863). Later, Bumba and Howard (1969) reported on regions that emerged repeatedly with rotation rates of 27 days and Gaizauskas et al. (1983) found that these complexes of activity were maintained for 3-6 solar rotations.
Multiple terms have been used to describe longitudinal clustering, and often the terminology is inconsistently used across the literature. Here we distinguish between related concepts: we use activity nests to describe clustering of AR emergence within a single hemisphere, while active longitudes refer to preferred emergence at similar longitudes in opposite hemispheres, implying symmetry across the equator (see Table 1). The term hot spot is used for regions that exhibit enhanced flare or sunspot productivity. Regardless of terminology, the longitudinal organization of solar magnetic activity provides a probe of non-axisymmetric components of the solar dynamo, which remain comparatively unexplored. Figure 1 depicts cartoon distributions of active regions: (left) moderate nesting similar to the Sun, (middle) moderate nesting with sunspots aligned along active longitudes, and (right) strong nesting as might be observed on a sun-like star, combined with well-defined active longitudes.
Quantitative definitions of activity nests were introduced by Castenmiller et al. (1986), who identified nests as at least three ARs emerging within four Carrington rotations and within 7.5∘ longitude and 5∘ latitude, finding that 30–40% of ARs between 1959–1964 participated in nests. The term activity nest is often used synonymously with active longitude. Some studies have described active longitudes as two preferred activity centers approximately 180∘ apart whose relative dominance alternates in time, a phenomenon termed as ’flip-flop’ (Berdyugina and Usoskin, 2003). Other investigations use the term active longitudes more generally to denote longitudinal concentrations or recurring zones of activity without requiring that the activity centers be 180∘ apart (Bogart, 1982; Bai, 1987, 2003; Henney and Harvey, 2002).
| Term | Historical Usage |
@@
| enhanced flaring; Bai (2003) | |
| Term | Proposed Usage |
| Activity Nests | One hemisphere only |
| Short-lived 6 months | |
| Long-lived 6 months | |
| Active Longitudes | Nests at similar longitudes |
| across the equator | |
| Hot Spots | Nests with flares |
The physical origin of longitudinal clustering remains uncertain. Proposed mechanisms include modulation by giant convective cells, magneto-Rossby waves, or MHD instabilities acting on toroidal flux bands near the tachocline (Zaqarashvili et al., 2010; Dikpati et al., 2018). Flux-emergence simulations further suggest that clustering may reflect subsurface connectivity among buoyant magnetic structures (Fan, 2009; Cheung et al., 2010). Persistent longitudinal preferences may therefore indicate non-axisymmetric dynamo modes or interactions between global rotation and convection. Despite decades of study, it remains unclear whether longitudinal clustering reflects a persistent dynamical structure of the solar dynamo or an emergent statistical property of flux emergence.
A central difficulty in identifying activity nests is that longitudinal organization depends on the assumed rotation frame. Most studies search for preferred longitudes in a frame rotating near the Carrington rate, implicitly assuming that longitudinal structures remain anchored in that frame (e.g., Bai, 1987, 2003). However, different rotation rates can either highlight or obscure persistent features. Bogart (1982) found rotation periods differing by up to two days between solar cycles, suggesting that preferred longitudes are not rigidly tied to a single rotation rate. Other studies reported migration of active longitudes consistent with differential rotation and cycle-dependent behavior (Berdyugina and Usoskin, 2003; Usoskin et al., 2005). In this work, we expand upon previous approaches by searching across a broad range of rotation rates, allowing nests and longitudinal clustering to be identified without presupposing their anchoring frame.
@@ 2 Data
For this study, we employ synoptic Carrington rotation maps of the vector magnetic field, , from the Helioseismic and Magnetic Imager (HMI; Scherrer et al. 2012; Schou et al. 2012) on board the Solar Dynamics Observatory (SDO). The synoptic maps provide full-Sun coverage of the photospheric magnetic field by assembling central meridian observations over a full Carrington rotation of 27.2753 days, enabling consistent analysis of the longitudinal distribution of solar activity. The maps can be accessed through JSOC in the data series hmi.synoptic_mr_polfil_720s. The dimensions are 3600 1440 with equal steps in longitude and sine latitude with values in the units of Mx cm-2.
In addition, we use of the Solar Photospheric Ephemeral Active Region (SPEAR) catalog http://hmi.stanford.edu/hminuggets/?p=3730, that is based on the Space-weather HMI Active Region Patch (SHARP) data product. SHARPs identify and track ARs in near-real time using HMI vector magnetograms (Bobra et al., 2014), and the SPEAR catalog provides a tabulated set of data by compiling AR properties such as total flux, latitude and longitude location, tilt angle, etc., when the SHARP regions are nearest to the central meridian. Each SHARP region is only found once in the catalog.
The monthly hemispheric sunspot numbers provided by the World Data Center for the Sunspot Index and Long-term Solar Observations (WDC-SILSO) are used to show the evolution of Solar Cycle 24 in the North and South hemispheres (WDC-SILSO, Royal Observatory of Belgium, 2025). Monthly values are calculated as averages of daily hemispheric counts and are distributed in the Version 2.0 SILSO data release, which provides a recalibrated and internally consistent extension of the international sunspot number series. The dataset, available since 1992, can be found here; https://www.sidc.be/SILSO/datafiles.
@@ 3 Methods
We use HMI Carrington Rotation maps of the radial vector magnetic field, , for visualization and comparison purposes only (not quantitative identification of nests), in order to see the times and locations of strong active region magnetic flux and compare the sunspot number with the radial magnetic flux time series from HMI. We plot the sum of the absolute value of between the latitudes of , in the top panel of Figures 2 and 3, keeping the hemispheres separate in order to note the northern and southern hemisphere patterns. We downsample the Carrington longitude from 3600 to 360 pixels and use a 2-dimensional Gaussian smoothing with a FWHM of 7∘ in longitude and 2 Carrington rotations in time. Note that the quantitative analysis to identify nests uses data in the SPEAR catalog, and does not have any smoothing. Also, since the Sun rotates differentially with higher latitudes rotating slower than those close to the equator, the Carrington rate is used to build the synoptic maps but is not necessarily indicative of the rate at which magnetic features rotate.
We used the SPEAR catalog to identify nests. ARs are recorded in the SPEAR catalog only once when they are closest to the central meridian. We remove long-lived ARs that return to the front side of the Sun for a second or third time. We remove returned ARs by first using a method developed by Yeates (2020) using flux transport models to estimate what a given AR rotating off the front side of the disk would look like upon its return. If the forward-modeled AR is a good match for the observed AR, then the observed AR NOAA number and SHARP number is listed as a returning sunspot in a separate file. The long-lived and returned ARs are then not allowed to be a member of any nest.
To identify a nest, each AR is used as a starting data point, then nearby latitudes (7.5∘), longitudes (5∘), and Carrington Rotations (4 Carrington Rotations) are searched to find other ARs that could be part of the nest given a certain rotation rate and nest definition. If a rotation rate is specified that is different from the Carrington rate of 424.3427 nHz, then the search for ARs in time updates the search locations to be longitude corresponding to the change in longitude per rotation that would correspond to that rotation rate. If an AR is identified in two nests, then those nests are merged. All frequencies in this paper are synodic, not sidereal. To convert to sidereal rates, add 31.6875 nHz to the synodic rate.
@@
In order to test the hemispheric asymmetry of ARs and nests, we first determine what fraction of the time an AR or nest in one hemisphere will have an AR or nest at the same longitude in the other hemisphere. We first plot the locations of all ARs from each hemisphere as a function of Carrington rotation and longitude and calculate how often they intersect, or occupy, the same longitude. This is the observed likelihood that an AR in one hemisphere will have an AR in a similar location in the other hemisphere. We then do the same for the identified nests. We report on how likely it is for a nest in one hemisphere to exist in the same longitude and time bin in the other hemisphere, simply from the intersection percentage. To test how likely the intersection of N-S nests are in a randomized distribution, the central longitude of each nest is randomized but the relative location of each AR within a given nest with respect to the central longitude is kept constant. This randomizes the nest locations in longitude but ensures the nest retains the same number of ARs. The time of each nest is not randomized for this test.
@@ 4 Results
The amount of nesting exhibited by the Sun depends on the definition used to describe what constitutes an activity nest. In this manuscript, we search for nests using the Castenmiller et al. (1986) definition with variable rotation rates.
@@ 4.1 Degree of Nesting
Using the criteria of Castenmiller et al. (1986), in which the nest critera is three ARs found within four CRs with 7.5∘ longitude and 5∘ latitude of each other, the SPEAR catalog is used to search through the AR locations to find clustering. When using the Carrington rate to search the catalog, 40 nests with 25% of ARs and 32% of flux are identified in the North and 30 nests with 23% of ARs and 35% of flux are found in the South, see the middle panel in Figures 2 and 3 that shows separate nests plotted as different symbols and colors. Note that latitude is not shown in the figures so that even if there is a large amount of flux in a location, a nest might not be found if the ARs are further away from each than 5∘ latitude.
If we used 6∘ in latitude to search for nests, the amount of flux in nests increases 2% in both the North and South hemispheres. If using 7∘ in latitude, it increases another 2-3%, continuing in this trend until it saturates at 15∘ latitude (the full width of the active latitudes) at the values of 43% (49%) flux found in nests in the N (S) hemisphere. The 5∘ in latitude represents a distance of 120 Mm, meaning ARs can be 120 Mm apart in latitude and still be in the same nest using this definition. Given that supergranules are 30 Mm in diameter, this would be a nest spanning four supergranules in latitude which is an appropriate, albeit large, distance for a nest in latitude. If we were searching for active longitudes, we would not limit the latitudinal extent at all.
The number of ARs and AR flux in nests increases when the data is searched using slightly prograde rates of rotation. The maximum amount of AR flux participating in nests is 48% (41%) in the South (North) hemisphere at rotational rates of 428.5 (426.7) nHz, slightly prograde with respect to the Carrington rate, see bottom panel in Figures 2 and Figure 3. These are considered short-lived nests since they live, on average, between 3.34 Carrington Rotations.
@@ 4.2 Nests at Different Rotation Rates
The percent of flux found in nests as a function of rotation rate is shown in Figure 4 for the Castenmiller et al. (1986) nest criteria, i.e. 3 ARs in 4 CRs, basically short-lived nests. The North and South hemisphere show significant flux in nests for rates near the Carrington rotation rate, i.e. synodic frequencies of 418-433 nHz synodic, or 27.7 -26.7 days. Other frequencies that contain significant flux in nests are found at 451-452 nHz and 409-413 nHz. These may be related to other processes such as inertial modes.
4.3 Nest Flux and Number During the Solar Cycle
The temporal evolution of magnetic flux contained within activity nests broadly follows the hemispheric solar cycle progression as traced by the monthly hemispheric sunspot number, see Figure 5. In both hemispheres, nest flux peaks near the epochs of maximum activity. Over the interval May 2010–December 2019, the total magnetic flux contained in nests is comparable between hemispheres, with G in the North and G in the South, corresponding to a North/South ratio of 0.91.
Despite the overall correlation with sunspot number, the relationship is not strictly proportional. For example, during mid-2012 in the Southern hemisphere the sunspot number exhibits a pronounced enhancement that is not accompanied by a corresponding increase in nest flux, indicating that elevated activity levels do not always manifest as strongly clustered emergence. The number of simultaneously present activity nests varies throughout the cycle, reaching maxima of five nests in the North and six in the South. In each hemisphere, the largest number of nests occurs close to the epoch of peak sunspot activity which is late 2011 for the North and late 2013 to early 2014 for the South, suggesting that nesting becomes most prevalent near cycle maximum.
@@ 4.4 Hemispheric Asymmetry
We examine the locations of all ARs that are assigned a NOAA number and also have an area greater than 50 Hemispheres. Separating the ARs into North and South hemispheres, we plot them as a function of Carrington Rotation and Carrington Longitude, see Figure 6, top panel. After dilating their locations by longitude and 3 Carrington rotations to provide some width, we count the percentage of ARs that intersect across the Northern and Southern hemisphere. The fraction of ARs that overlap in North and South bins with this dilation is 58%. This can be interpreted as follows: given an AR in one hemisphere, there is a 58% chance that an AR will be located within 9∘ longitude and 3 Carrington rotations in the opposite hemisphere.
The same is done for the nest population of ARs, see the middle panel of Figure 6, using the Castenmiller criteria with a Carrington rotation rate. Of all the longitude-time bins occupied by an AR in a nest in one hemisphere, only 7.8% have an AR in a nest in the opposite hemisphere within 3 CR in time and in longitude. This means that when nests form in one hemisphere, they are overwhelmingly absent at the corresponding longitude and time in the opposite hemisphere. In contrast to the broader active region population, which shows substantial co-occupancy, the organization of nests is characterized by strong hemispheric antisymmetry. Identifying nests using a slightly prograde rotation rate (the rates shown in Figures 2 and 3 when the maximum amount of flux is found in nests) provides a North-South ARs in nests intersection of 11.9%.
We determine the significance of the nest asymmetry by randomizing the central longitude of the nest locations. Meaning, we characterize the asymmetry by randomly shifting the average longitude of each nest while keeping the number of ARs in the nest constant. We do not randomize the latitude or time of the nests. The 10,000 trials show that the observed nest asymmetry is greater than 92% of the randomized trials, see Figure 6 lower panel. If using the slightly prograde rotation rates to find nests, the nest asymmetry is greater than 83.4% of trials.